Two trains leave the station at the same time. Train A travels at a constant speed of 16 m/s. Train B starts at 8.0 m/s but accelerates constantly at 1.0 m/s^2. After 10 seconds, which train has the greater speed? Explain.
1
Expert's answer
2020-01-03T09:19:35-0500

The speed of a train A


"v_A=16\:\rm m/s"

The speed of a train B


"v_B=v_0+at=8.0+1.0\times 10=18\rm \: m/s"

Hence


"v_B>v_A"